OpenStudy (anonymous):

Solve the following system of linear equations. 3x + 2y = 10 2x + 3y = 15/2

OpenStudy (anonymous):

3x + 2y = 10 2x + 3y = 15/2 Add two equations 5x+5y=35/2 x+y=7/2 2x+2y=7 3x+2y=10 2x+2y=7 x=3 Plug in x to get y.